Consider the R2−R function f defined by f(x,y) = x−2y. Prove from first principles that lim(x,y)→(2,1) f(x,y)=0.


1
Expert's answer
2021-05-14T03:35:02-0400

Ans:-

Given that "f(x,y) = x\u22122y" which belongs to R2−R function


"\Rightarrow lim_{(x,y)→(2,1)} f(x,y)"


applying first principal

"\Rightarrow lim_{h\to 0}\dfrac{f(x+h,y+h)-f(x)}{h}\\
\Rightarrow lim_{h\to 0}\dfrac{(x+h)-2(y+h)-(x-2y)}{h}\\
\Rightarrow =0"

Hence "lim_{(x,y)\u2192(2,1)} f(x,y)=0"